# Treat Visual Tokens as Text? But Your MLLM Only Needs Fewer Efforts to See

### v2

@@ Abstract
By treating visual tokens from visual encoders as text tokens, Multimodal Large Language Models (MLLMs) have achieved remarkable progress across diverse visual understanding tasks, leveraging the robust architectures of Large Language Models (LLMs). However, as token counts grow, the quadratic scaling of computation in LLMs introduces a significant efficiency bottleneck, impeding further scalability. Although recent approaches have explored pruning visual tokens or employing lighter LLM architectures, the computational overhead from an increasing number of visual tokens remains a substantial challenge.
In this study, we investigate the redundancy in visual computation at both the parameter and computational pattern levels within LLaVA, a representative MLLM, and introduce a suite of streamlined strategies to enhance efficiency. These include neighbor-aware visual token attention, pruning of inactive visual attention heads, and selective layer dropping for visual computations. By implementing these strategies in LLaVA, we achieve a reduction in computational demands of $88\%$ while maintaining model performance across key benchmarks. Additionally, we validate the existence of visual computational redundancy in other MLLMs, such as Qwen2-VL-7B and InternVL-2.0-4B/8B/26B. These results present a novel pathway for MLLMs to handle dense visual tokens with minimal computational costs. Code and model checkpoints will be released to support further research.
@@ 1 Introduction
In our work, we seek to fully leverage visual redundancy to enhance the efficiency of MLLMs. Rather than pruning input visual tokens to reduce redundant visual representations, we explore redundancy within visual computation patterns in MLLMs. Using LLaVA as a case study, we conduct a detailed analysis of redundancies in both visual attention and representation computations across all layers. Based on these insights, we introduce a set of simple yet effective strategies to significantly reduce the computational burden from visual information: neighbor-aware visual attention, non-active visual attention head dropping, sparse projection in the FFN, and lazy layer dropping for visual processing. As shown in Fig. 1, most parameters are dropped during visual computation while remaining active in text computation, thereby maintaining the strong understanding capability of the LLM while efficiently reducing the computational overhead introduced by the large number of visual tokens. By integrating these strategies into LLaVA, we reduce visual computational overhead by up to $88\%$ while retaining nearly the same performance as the original model. Across multiple benchmarks, our method achieves state-of-the-art performance compared to various baselines. Furthermore, these findings are not unique to the LLaVA-based model; similar visual computational redundancies are also present in other MLLMs, including Qwen2-VL-7B and InternVL-2.0-4B/8B/26B. These results highlight a promising direction for future research to enhance MLLM efficiency.
@@
    We introduce three strategies for visual computation pruning: neighbor-aware visual attention, non-active visual attention head dropping, sparse visual projection in the FFN, and lazy layer dropping for visual processing. These strategies effectively reduce computational costs at the inherent computation pattern level, alleviating the increasing visual computation overhead associated with a large number of visual tokens.
@@ 3.1 Redundancy of visual computation in MLLMs
Despite this significant computational load, much of the visual processing in MLLMs is redundant due to the inherent spatial sparsity in visual data. To illustrate this, we randomly select $10$ visual tokens from three example images and visualize their attention weights with respect to other vision tokens in shown images, i.e., $Q_{v}K_{v}^{T}$ in Eq. 1. More specifically. we first compute the average attention weights across all attention heads and then rearrange these weights for each selected token based on its spatial distance from other tokens. This enables us to investigate the relationship between token spatial distance and the redundancy in visual attention computation. The spatial distance between visual token $i$ and $j$ is defined as follows,
@@ 4.1 Experiment setup
Implementations. In our experiments, we apply individual and combined pruning strategies to reduce visual computation redundancy in the model. We evaluate our approach in two settings: training-based and training-free. In the training-based setting, we use the same data and training protocol as the public repository of LLaVA-1.5 models, training from scratch with reduced visual computation redundancy. The number of iterations remains the same to allow a fair comparison with the original model. In the training-free setting, we directly apply different pruning strategies to studied models, i.e., Qwen2-VL-7B, and InternVL-2.0, without any fine-tuning. All experiments were conducted on a single node with 8 A100-80G GPUs.
@@ 4.4 Computational redundancy beyond LLaVA
The visual computational redundancy is not unique to LLaVA. To validate the broader applicability of our pruning strategy, we applied it to additional MLLMs, including Qwen2-VL-7B and InternVL-2.0, without fine-tuning due to limited training data. We evaluated performance on GQA and POPE benchmarks, adjusting pruning granularity to match original model performance with minimal FLOPs. As shown in Fig. 7, we observe that, even without fine-tuning, pruning visual computations in these models at suitable ratios does not compromise their performance. Furthermore, larger MLLMs appear to accommodate higher pruning ratios, as evidenced by the results from pruning InternVL-2.0 at various model scales.

### v3

@@ Abstract
By treating visual tokens from visual encoders as text tokens, Multimodal Large Language Models (MLLMs) have achieved remarkable progress across diverse visual understanding tasks, leveraging the robust architectures of Large Language Models (LLMs). However, as token counts grow, the quadratic scaling of computation in LLMs introduces a significant efficiency bottleneck, impeding further scalability. Although recent approaches have explored pruning visual tokens or employing lighter LLM architectures, the computational overhead from an increasing number of visual tokens remains a substantial challenge. It also raises challenges to decide important visual tokens for different sample, which introduces another computation overhead.
In this study, we investigate the redundancy in visual computation at both the parameter and computational pattern levels within LLaVA, a representative MLLM, and introduce a suite of streamlined strategies to enhance efficiency. These include neighbor-aware visual token attention, pruning of inactive visual attention heads, and selective layer dropping for visual computations. By implementing these strategies in LLaVA, we achieve a reduction in computational demands of $88\%$ while maintaining model performance across key benchmarks. Additionally, we validate the existence of visual computational redundancy in other MLLMs, such as Qwen2-VL-7B and InternVL-2.0-4B/8B/26B. These results present a novel pathway for MLLMs to handle dense visual tokens with minimal computational costs. Our pruning method only requires one time to discover the important visual-related computation and is sample-agnostic. This indicates that, You Only need Pune Once (YOPO) to accelerate your MLLMs. Code and model checkpoints are available at https://github.com/ZhangAIPI/YOPO_MLLM_Pruning.
@@ 1 Introduction
In our work, we seek to fully leverage visual redundancy to enhance the efficiency of MLLMs. Rather than pruning input visual tokens to reduce redundant visual representations, we explore redundancy within visual computation patterns in MLLMs. Using LLaVA as a case study, we conduct a detailed analysis of redundancies in both visual attention and representation computations across all layers. Based on these insights, we introduce a set of simple yet effective strategies to significantly reduce the computational burden from visual information: neighbor-aware visual attention, non-active visual attention head dropping, sparse projection in the FFN, and lazy layer dropping for visual processing. As shown in Fig. 1, most parameters are dropped during visual computation while remaining active in text computation, thereby maintaining the strong understanding capability of the LLM while efficiently reducing the computational overhead introduced by the large number of visual tokens. By integrating these strategies into LLaVA, we reduce visual computational overhead by up to $88\%$ while retaining nearly the same performance as the original model. Across multiple benchmarks, our method achieves state-of-the-art performance compared to various baselines. Furthermore, these findings are not unique to the LLaVA-based model; similar visual computational redundancies are also present in other MLLMs, including Qwen2-VL-7B and InternVL-2.0-4B/8B/26B. While token pruning-based methods require dynamic identification of important visual tokens on a per-request basis, our method directly prunes the computation pattern in MLLM from the outset of model development. This means that You Only Need to Prune Once (YOPO), which effectively reduces visual computation redundancy and accelerates MLLM performance across all subsequent use cases. These results highlight a promising direction for future research to enhance MLLM efficiency.
@@
    We introduce three strategies for visual computation pruning: neighbor-aware visual attention, non-active visual attention head dropping, sparse visual projection in the FFN, and lazy layer dropping for visual processing. These strategies effectively reduce computational costs at the inherent computation pattern level, alleviating the increasing visual computation overhead associated with a large number of visual tokens. Notably, our method prunes the MLLM only once, and it can be deployed for all subsequent computation requests.
@@ 3.1 Redundancy of visual computation in MLLMs
Despite this significant computational load, much of the visual processing in MLLMs is redundant due to the inherent spatial sparsity in visual data. To illustrate this, we randomly select $10$ visual tokens from three example images and visualize their attention weights with respect to other vision tokens in shown images, i.e., $Q_{v}K_{v}^{T}$ in Eq. 1. More specifically, we first compute the average attention weights across all attention heads and then rearrange these weights for each selected token based on its spatial distance from other tokens. This enables us to investigate the relationship between token spatial distance and the redundancy in visual attention computation. The spatial distance between visual token $i$ and $j$ is defined as follows,
@@ 4.1 Experiment setup
Implementations. In our experiments, we apply individual and combined pruning strategies to reduce visual computation redundancy in the model. We evaluate our approach in two settings: training-based and training-free. In the training-based setting, we use the same data and training protocol as the public repository of LLaVA-1.5 models, training from scratch with reduced visual computation redundancy. The number of iterations remains the same to allow a fair comparison with the original model. In the training-free setting, we directly apply different pruning strategies to studied models, i.e., Qwen2-VL-7B, and InternVL-2.0, without any fine-tuning. All experiments were conducted on a single node with 8 A100-80G GPUs, with detailed results provided in Appendix A.
@@ 4.4 Computational redundancy beyond LLaVA
The visual computational redundancy is not unique to LLaVA. To validate the broader applicability of our pruning strategy, we applied it to additional MLLMs, including Qwen2-VL-7B and InternVL-2.0, without fine-tuning due to limited training data. We evaluated performance on GQA and POPE benchmarks, adjusting pruning granularity to match original model performance with minimal FLOPs. As shown in Fig. 7, we observe that, even without fine-tuning, pruning visual computations in these models at suitable ratios does not compromise their performance. Furthermore, larger MLLMs appear to accommodate higher pruning ratios, as evidenced by the results from pruning InternVL-2.0 at various model scales. More details can be found in Appendix C.
@@ 5 Conclusion
Appendix A Experiment details
For the LLaVA pruning with training. We employed training data and settings identical to those used in LLaVA-1.5. All experiments were conducted on a single node equipped with $8$ A100-80G GPUs. The pre-training phase took $4$ hours, while instruction fine-tuning took 20 hours.
For the MLLMs pruning without training. For the experiments of pruning models without the training for performance restoration, we use one A100-80GB GPU.
Pruning details for our models. In Tab. 2, we present the results of the LLaVA pruned by the combination of our strategies. Here, we present the pruning details in Tab. 6.
Appendix B More results on LLaVA-1.5-7B
We provide a detailed evaluation of pruning LLaVA-1.5-7B using various strategies, including neighbor-aware attention computation (Tab. 7), dropping inactive attention heads (Tab. 8), and layer dropping (Tab. 9). The results show that a radius of 5 achieves efficient attention computation with strong performance. Additionally, dropping 12 attention heads and the last 12 layers during visual computation allows for significant acceleration without compromising performance.
Appendix C More results on Qwen2-VL-8B and InternVL-2.0-4B/8B/26B
We present more results of pruning the Qwen-2-VL-8B and InternVL-2.0-4B/8B/26B without the fine-tuning under different pruning strategies and granularity.
Pruning details. For Qwen-2-VL-8B, we respectively use the non-active attention head dropping and neighbor-aware attention computation to reduce the FLOPs to 75%. For InternVL-2.0-4B/8B, we set the radius for attention computation as $4$ to reduce the FLOPs to $75\%$, and integrate the non-active attention head dropping with $24$ attention heads maintaining to further reduce to 66%, then discard the visual-related computation after $24$ layers to reduce to 50%, while changing the layer from $24$ to $16$ reduces the FLOPs to 44%. For InternVL-2.0-8B, we further reduce the last layer for visual computation to $12$ to reduce the FLOPs to 33%. For InternVL-2.0-26B, by setting the radius for visual attention computation as $4$, maintaining $24$ active visual attention heads, and discarding the visual-related computation after the $36$-th layer, we can reduce 50% FLOPs, while changing the discarding layer to $24$ and $19$ can further reduce to 33% and 25% FLOPs, respectively. The results are shown in Fig. 9 and Fig. 10
Table 6: Summary of LLaVA-ours models with corresponding pruning parameters. We denote # Heads as the number of attention heads retained for the visual attention computation, and Radius as the radius of neighbor visual tokens for attention computation. The # Layers is denoted as the number of last layers to drop for the visual-related computation. The # Neurons is denoted as the ratio of neurons retained at the FFN layer.
Model FLOPs # Heads Radius # Layers # Neurons
LLaVA-7B-ours 1.91 24 5 16 50%
0.92 16 25%
LLaVA-13B-ours 3.72 30 5 20 50%
1.79 20 25%
Table 7: Results of applying the neighbor-aware attention mechanism as a visual pruning strategy to the LLaVA-1.5-7B model. The ’radius’ parameter defines the scope of neighboring tokens considered in visual token computation.
Radius GQA VQAv2 SQA TextVQA Avg
1 61.8 78.6 69.9 56.9 66.8
3 61.9 78.8 69.4 57.3 66.9
5 62.2 78.9 69.0 57.9 67.0
7 61.9 79.0 69.1 58.6 67.1
9 62.2 79.0 68.7 58.8 67.2
11 60.6 78.2 69.2 57.6 66.4
Table 8: Results of dropping non-active attention heads as a visual pruning strategy to prune the LLaVA-1.5-7B model. ’# Heads’ indicates the number of dropped non-active attention heads per layer used in visual token computation.
# Heads GQA VQAv2 SQA TextVQA Avg
4 62.1 78.9 68.0 58.3 66.8
12 61.9 78.6 68.3 57.3 66.5
20 61.6 78.1 67.7 57.2 66.1
28 60.8 77.1 67.7 56.0 65.4
Table 9: Results of applying the layer-dropping strategy to prune the LLaVA-1.5-13B model. # layers indicates the number of last several layers to drop for visual-related computation.
# Layers GQA VQAv2 SQA TextVQA Avg
4 62.4 79.1 68.7 57.8 67.0
8 62.2 79.1 68.7 58.3 67.0
12 62.2 79.0 68.9 57.5 67.0
16 62.1 78.7 68.4 57.0 66.6
20 61.4 78.0 68.0 55.3 65.7
24 59.3 75.4 68.1 54.0 64.2
28 54.5 71.1 68.4 51.6 61.4
Figure 9: Evaluation results of Qwen2-VL-8B under different pruning strategies. We denote D.H. as dropping non-active visual attention heads during the visual-related computation, and N.A. as the neighbor-aware visual attention computation.
Figure 10: Evaluation results of InternVL-2.0-4B/8B/26B under different pruning ratios. The ratio of $1$ indicates the original model without pruning.
It can be shown that even without any fine-tuning, our proposed pruning strategies can efficiently reduce the required FLOPs with remaining similar performance compared with the original model. It should be noticed that from the results of pruning InternVL-2.0 models, we can see that a larger model usually has a larger computation redundancy. For example, while the use of $44\%$ FLOPs causes a large performance drop for InternVL-2.0-8B model, the InternVL-2.0-26B model still has a comparable performance compared with the original one. Also, for a given pruning ratio, we should carefully select the pruning strategy. From the results of Qwen2-VL-8B, the use of neighbor-aware attention computation could cause a larger performance drop than the use of non-active attention head dropping. It remains further work to study how to select a better combination of pruning strategies to get a good performance with improved efficiency.
Appendix D Demo application
We also present three real-world examples to show the performance of the model (LLaVA-1.5-13B) before and after pruning. The results are depicted in Fig. 11, Fig. 12, and Fig. 13. We highlight the expected keywords in violet. Even with only 12% of FLOPs allocated to visual-related computation, MLLM demonstrates the ability to accurately capture semantic information. This highlights the significant potential of MLLM visual computation pruning in accelerating real-world applications.
Figure 11: Demo application result #1.
Figure 12: Demo application result #2.
Figure 13: Demo application result #3.